A cone has a volume of 108(pi) cubic inches and a height of 4 inches. What is its radius?
Answer by mananth(16949)

Volume of a cone = 1/3 * pi*r^2*h
108*pi=1/3 *pi*r^2 * 4
108*pi/pi*4=1/3r^2
108/4= 1/3 *r^2
27=1/3*pi*r^2
81=r^2
sqrt81= r
9= r
Radius = 9 inches